Suppose that the annual rate of return for a common biotechnology stock is normally distributed with a mean of 4% and a standard deviation of 4% . Find the probability that the one-year return of this stock will be positive. Round your answer to at least four decimal places.
Answer by stanbon(48568)

Suppose that the annual rate of return for a common biotechnology stock is normally distributed with a mean of 4% and a standard deviation of 4% . Find the probability that the one-year return of this stock will be positive. Round your answer to at least four decimal places.
---------------
Find the z-score of zero:
z(0) = (0-0.04)/(0.04) = -0.04/0.04 = -1
-------------
P(p>0)= P(z>-1) = 0.8413...
Cheers,
Stan H.